The table shows a short excerpt from the car weight and mileage data file on the text CD. That file lists several 2004 mode**DROP DOWN OPTIONS FOR B: FEWER OR MORE
***DROP DOWN OPTIONS FOR C: MILEGE or WEIGHT // MILEGE or WEIGHT
Answer #1

we have the regression model do poedict the mileage Y =47.905 - 0.0056 = ý = 47.905339-0.005619*7 (without rounding) from exc